find three consecutive even integers whose sum is 114
Answer by umail08(87)

x = first consecutive even integer
x + 2 = second consecutive even integer
x+4 = third consecutive even integer
x + x +2 + x + 4 = 114
3x + 6 = 114
3x + 6 – 6 = 114 – 6
3x = 108
x = 36
x + 2 = 38
x + 4 = 40